# An All-digital 8.6-nJ/Frame 65-nm Tsetlin Machine Image Classification Accelerator

### v2

@@ Abstract
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Current methods for accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3]. CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources. As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3]. The complexity and energy consumption can be reduced by compromising the data precision [4]. Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [10] is an ML algorithm founded on propositional logic. Parallel conjunctive clauses are defined by independent learning automata, and each clause represents a logical sub-pattern. The TM features a single-layer structure with highly interpretable outputs, making it well-suited for application areas such as biomedical, cybernetics, IoT, and law. Furthermore, because its primary operations are of Boolean type, TM solutions are hardware-efficient. The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training. This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [11]. Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [12].
@@
Until now, only a single ASIC based on the TM has been reported [11], which operated on a simple dataset. We wanted to make an an accelerator that operated on a well-known dataset (MNIST) and to demonstrate that an all-digital TM-based ASIC could achieve energy-efficiency at the same level as analog in-memory-computing (IMC) solutions [20, 21, 9].
The ASIC’s accelerator core is built on a modified version of the ConvCoTM inference core included in the FPGA solution in [12], which supported full on-device training. However, because of the FPGA it could not achieve ultra-low-power (ULP) operation. Compared to [12], several enhancements are made in the inference core of the ASIC, as described in section IV. In particular, there is an updated clause logic circuit with a novel feedback that reduces switching activity. There is also new functionality for loading a pre-trained model. Most important is the optimization of the inference part to achieve ULP operation.
@@
  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of the clause’s combinational output.
@@
The remainder of the paper is organized as follows: Section II summarizes related work. Section III covers TM and CoTM background, including convolution operation. Section IV details the ASIC’s architecture and building blocks. Measurement results are presented in Section V, followed by a discussion in Section VI before we conclude the paper.
@@ II Related Work
In [7] a digital 10 nm FinFET CMOS solution is reported. Utilizing a BNN algorithm, it achieves a peak energy efficiency of 617 TOPS/W (Tera operations per second per watt). It has a power consumption of 5.6 mW and obtains an accuracy of 86% when operating on the CIFAR-10 dataset.
@@
In [9] a charge-domain IMC ternary neural network (TNN) chip is described. The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model. The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W. This corresponds to an EPC of 0.18 $\mu$J. The test accuracy achieved is 97.1%.
@@
Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption. In [21] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology. It achieves a test accuracy on MNIST of 95.35%, with a power consumption of 0.517 mW and an EPC of 12.92 nJ. Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [24]. The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%. In [25], simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described. On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
@@
In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC as described in the Introduction section. It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes full on-device training. For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134k samples per second, and an EPC of 13.3 $\mu$J. The source code for this design is available at [28].
@@ III-A General TM Concepts
The input feature vector to a TM consists of $o$ propositional Boolean variables, $x_{u}\in\{0,1\}^{o}$, where $u\in\{0,\ldots,o-1\}$. A new input vector named the literals, L, is generated by appending the negated variables to the input, as shown in Eq. (1). Thus, there are in total $2o$ literals.
@@
A TA of two-action type [34] will either include or exclude a literal in a conjunctive clause, depending on its state. One team of TAs is applied per clause, and the different TA actions are obtained during the training process, which involves feedback mechanisms to each TA [10]. For an inference-only solution, as reported in this paper, only the TA action signals from the trained model are needed, not the complete TAs.
@@
where $l_{k}$ is the literal with index k. $k$ belongs to $I_{j}\subseteq\{0,\dots,2o-1\}$, where $I_{j}$ is the set of indices of the literals in clause $j$ for which the corresponding TAs take the action include.
@@ III-B CoTM
The clause weights in Eq. (3) can have both positive and negative polarity, and obtain their values during training. No multiplications are needed to implement Eq. (3), only additions, as a clause can output either 0 or 1. The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq. (4).
@@ III-C CoTM and Convolution
In Eq. (5), $U$ is the number of bits used for thermometer encoding [35] of the value of a single pixel (one-hot encoding can also be used). Such encoding is required because a TM operates on Boolean variables, not on binary numbers. The term $(Y-W_{Y})+(X-W_{Y})$ in Eq. (5) represents the number of bits that encode the patch’s position [19, 12] in the $Y$ and $X$ directions respectively. Also here we apply thermometer encoding. As shown in Eq. (1), the ConvCoTM also takes as input the negated versions of the features to form the literals per patch.
@@ III-D Datasets and Booleanization
The widely used MNIST [14] dataset consists of handwritten digits. FMNIST [15] is more challenging than MNIST and includes images from a Zalando catalog, such as t-shirts, sandals, and shoes. KMNIST [16] consists of cursive Japanese characters, and is also more demanding than MNIST. Each of these datasets includes images of $28\times 28$ pixels organized in 60k training samples and 10k test samples. There are 10 classes within each dataset. As the images are of single-channel (greyscale) type, $Z=1$.
The pixel values range from 0 to 255. For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [13]. For FMNIST and KMNIST, an adaptive Gaussian thresholding procedure [13] is applied. We use one bit per pixel, i.e., $U=1$ for all three datasets. The convolution window has dimensions $W_{X}=W_{Y}=10$, and a step size of one is used in both directions. For each patch, we have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref. Eq. (5). The number of literals for each patch is therefore 272.
@@ III-E TM and Convolution - Inference
During inference, each clause in a ConvCoTM is evaluated $B$ times per image. After finishing the patch generation, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [13]. The clause output is given by Eq. (6) which we call the sequential OR-function. Here $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the patch generation:
@@ IV-A Interface to the System Processor
The 8-bit parallel data interface towards the external system processor, is inspired by the ARM AXI Stream interface [36]. It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode. When a classification operation is finished, the accelerator outputs an interrupt. In addition, it outputs a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
@@ IV-C Image Data Buffer and Patch Generation
The patch generation module consists of a register with 10 rows, each with 28 DFFs. At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates ($0,0$).
During the patch generation, the window is moved to the right with increments of one. As described in Subsection III-C, the position of the window is thermometer encoded. There are 19 different $x$-coordinates for the window, requiring 18 bits. When the window has been evaluated at the rightmost $x$-position ($0,18$), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost register row. The convolution window then starts again at the leftmost position, now with patch coordinates ($1,0$). This procedure is repeated until all 28 datarows of the image have been processed. There are 19 different positions also for the $y$-coordinate, and during a complete patch generation phase, $19\times 19=361$ patches are generated.
@@ IV-D Clause Pool
In this accelerator ASIC we apply a single convolution window. If increased classification rate is wanted, one can use several convolution windows that operate in parallel. In this case, the combinational clause logic that generates the signal $c_{j}^{b}$ in  Fig. 2, would have to be replicated for each convolution window.
@@ IV-E Inference
The inference operation starts with loading an image into the data buffer, followed by sequential generation of the patches, as described in Subsection IV-C. When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq. (3).
All ten class sums are generated in parallel, by adders configured in a tree-structure and utilizing a three-stage pipeline. Finally, an argmax module [11] selects the class with the highest class sum, as the predicted class, $\hat{y}$. Fig. 3 shows a simplified state diagram of the accelerator.
@@
When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C. This increases the throughput, as image samples can be processed every 372’th clock cycle. Any timing overhead in the system processor will add to the total latency.
@@ IV-F Clock Domains
V Implementation and Measurement Results
The ConvCoTM accelerator was designed in VHDL and implemented in a 65 nm low-leakage CMOS technology from UMC. A JLCC44 ceramic package housed the chip. Fig. 4 shows a photograph and a layout plot of the ASIC. The accelerator’s core area is approximately 2.7 mm2. The VHDL source code for the ASIC implementation (used for initial FPGA verification) is available at a GitHub repository [37].
A printed circuit board (PCB) was developed for testing and characterization of the ASIC. The ASIC PCB was connected to a Digilent Zybo Z7-20 development board, equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig. 5. One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC PCB. The processor also read back accelerator status signals, and the predicted class per sample. To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage. The accelerator core operated from 0.82 V to 1.2 V.
@@ V Implementation and Measurement Results
TABLE I: The ConvCoTM accelerator ASIC’s main characteristics and performance

### v3

@@ Abstract
publicationid: pubid: Copyright 2025 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission.
@@ I Introduction
Current methods for accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3]. CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources. As such, they require a large number of micro-architectural operations for each classification and have significant energy footprints [3]. The complexity and energy consumption can be reduced by compromising the data precision [4]. Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [10] is an ML algorithm founded on propositional logic. Parallel conjunctive clauses are defined by independent learning automata, and each clause represents a logical sub-pattern. The TM features a single-layer structure with highly interpretable outputs, making it well-suited for application areas such as biomedical, cybernetics, IoT, and law. Furthermore, because its primary operations are of Boolean type, TM solutions are HW-friendly. The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training. This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [11]. Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [12].
@@
Until now, only a single ASIC based on the TM has been reported [11], and it operated on a simple dataset. We wanted to make an an accelerator that operated on a well-known dataset (MNIST) and to demonstrate that an all-digital TM-based ASIC could achieve energy-efficiency at the same level as analog in-memory-computing (IMC) solutions [20, 21, 9].
The ConvCoTM FPGA solution in [12] demonstrated inference and full-on device training. However, ultra-low-power (ULP) operation could not be achieved with the FPGA. As full on-device TM-based training had been proved feasible in [12], we wanted to design an inference-only ASIC solution, optimized for low-power operation.
The ASIC accelerator’s inference core is built on modules included in [12], with several enhancements as described in section IV. In particular, there is an updated clause logic circuit with a novel feedback that reduces switching activity. There is also new functionality for loading a pre-trained model. Most important is the optimization of the inference part to achieve ULP operation.
@@
  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of a clause’s combinational output.
@@
The remainder of the paper is organized as follows: Section II summarizes related work. Section III covers TM and CoTM background, including convolution operation. Section IV details the ASIC’s architecture and building blocks, and measurement results are presented in Section V. In Section VI we discuss possible extensions and modifications of the design. The ConvCoTM accelerator’s performance is discussed in Section VII, and Section VIII concludes the paper.
@@ II Related Work
In [7] a digital 10 nm FinFET CMOS solution is reported. Utilizing a BNN algorithm, it achieves a peak energy efficiency of 617 Tera operations per second per watt (TOPS/W). It has a power consumption of 5.6 mW and obtains an accuracy of 86% when operating on the CIFAR-10 dataset.
@@
In [9] a charge-domain IMC ternary neural network (TNN) chip is described. The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model. The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W, which corresponds to an EPC of 0.18 $\mu$J. The test accuracy achieved is 97.1%.
@@
Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption. In [21] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology. It achieves a test accuracy on MNIST of 95.35%, with a power consumption of 0.517 mW and an EPC of 12.92 nJ. Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [24]. The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%. In [25], simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described. On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC. A time-domain IMC mixed-signal SNN is reported in [26]. It is implemented in a 65 nm CMOS technology and has an energy efficiency of 701.7 TOPS/W. On the CIFAR-10 dataset it achieves a test accuracy of 91.3%.
In [27] an analog IMC accelerator suited for CNNs as well as for Transformers is presented. It achieves 818–4094 TOPS/W, and CIFAR-10 test accuracies of 91.7% and 95.8% for CNN (ResNet20) and Transformer (ViT-S) respectively. The technology used is 65 nm CMOS.
@@
In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC, as described in the Introduction section. It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes full on-device training. For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134 k samples per second, and an EPC of 13.3 $\mu$J. The source code for this design is available at [30].
The FPGA-based accelerator in [31] is a flexible solution that supports on-device training for the vanilla TM and the CoTM. It requires less HW resources and provides greater flexibility compared to [12], at the expense of higher latency.
@@ III-A General TM Concepts
The input feature vector, $X$, to a TM consists of $o$ Boolean variables, $X=[x_{0},x_{1},\dots,x_{o-1}]$. A new input vector named the literals, L, is generated by appending the negated variables to the input, as shown in Eq. (1). Thus, there are in total $2o$ single-bit literals.
@@
A TA of two-action type [37] will either include or exclude a literal in a conjunctive clause, depending on its state. Fig. 1 shows the structure of a single TA with $2N$ states. In HW, a TA is typically implemented as a binary up/down counter, and the inverted version of its most significant bit (MSB) is used as the TA action signal (active high).
Figure 1: A Tsetlin Automaton (TA) for two-action environments.
One team of TAs is applied per clause, and the different TA actions are obtained during the training process, which involves feedback mechanisms to each TA [10]. For an inference-only solution, as reported in this paper, only the TA action signals from the trained model are needed, not the complete TAs.
@@
where $l_{k}$ is the literal with index k. $k$ belongs to the set $I_{j}\subseteq\{0,\dots,2o-1\}$, which is obtained from training, or from a pre-trained model. $I_{j}$ consists of the indices of the literals in clause $j$ for which the corresponding TAs take the action include.
@@ III-B CoTM
The clause weights in Eq. (3) can have both positive and negative polarity, and obtain their values during training. No multiplications are needed to implement Eq. (3), as a clause’s output is either 0 or 1. The predicted class, $\hat{y}$, is determined by the argmax operator:
@@ III-C CoTM and Convolution
In Eq. (5), $N_{F}$ corresponds to $o$, which is the number of Boolean variables which are input to the TM, as described in Subsection III-A. $U$ is the number of bits used for thermometer encoding [38] of the value of a single pixel (one-hot encoding can also be used). Such encoding is required because a TM operates on Boolean variables, not on binary numbers. The term $(Y-W_{Y})+(X-W_{Y})$ in Eq. (5) represents the number of bits that encode the patch’s position [19, 12] in the $Y$ and $X$ directions respectively. Also here we apply thermometer encoding. As shown in Eq. (1), the ConvCoTM also takes as input the negated versions of the features to form the literals per patch.
@@ III-D Datasets and Booleanization
The widely used MNIST dataset consists of handwritten digits. FMNIST is more challenging than MNIST and includes images from a Zalando catalog, such as t-shirts, sandals, and shoes. KMNIST consists of cursive Japanese characters, and is also more demanding than MNIST. Each of these datasets includes images of $28\times 28$ pixels organized in 60k training samples and 10k test samples. There are 10 classes within each dataset. As the images are of single-channel (greyscale) type, $Z=1$.
The pixel values range from 0 to 255. For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [13]. For FMNIST and KMNIST, an adaptive Gaussian thresholding procedure [13] is applied. We use one bit per pixel, i.e., $U=1$ for all three datasets. The convolution window has dimensions $W_{X}=W_{Y}=10$, and a step size of one is used in both directions. For each patch, we therefore have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref. Eq. (5). Thus, there are 272 literals per patch.
@@ III-E TM and Convolution - Inference
During inference, each clause in a ConvCoTM is evaluated $B$ times per image. After finishing the patch generation, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [13]. The clause output is given by Eq. (6) which we denote as the sequential OR-function. Here $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the patch generation:
@@ IV-A Interface to the System Processor
The 8-bit parallel data interface towards the external system processor, is inspired by the ARM AXI Stream interface [39]. It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode. When a classification operation is finished, the accelerator outputs an interrupt, and a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
@@ IV-C Image Data Buffer and Patch Generation
The patch generation module consists of a register with 10 rows, each with 28 DFFs. At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates ($0,0$), see Fig 3.
Figure 3: Register structure for window sliding. Each square represents a DFF. The first 10 rows of a booleanized image have been loaded and the window is in the start position (0,0).
During the patch generation, the window is moved to the right with increments of one. As described in Subsection III-C, the position of the window is thermometer encoded. There are 19 different $x$-coordinates for the window, requiring 18 bits. When the window has been evaluated at the rightmost $x$-position ($0,18$), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost register row. The convolution window then starts again at the leftmost position, now with patch coordinates ($1,0$). This procedure is repeated until all 28 datarows of the image have been processed. There are 19 different positions also for the $y$-coordinate, and during a complete patch generation phase, $19\times 19=361$ patches are generated. The $x$- and $y$-positions are thermometer encoded, each with 18 bits as shown in Table I.
TABLE I: Thermometer position encoding of $10\times 10$ convolution window in a $28\times 28$ image. $x$ at position 0 means leftmost column, while $y$ at position 0 means uppermost row.
$x$ or $y$ position Thermometer encoded value (18 bits)
0 000000000000000000
1 000000000000000001
$\cdots$ $\cdots$
17 011111111111111111
18 111111111111111111
@@ IV-D Clause Pool
In this accelerator ASIC we apply a single convolution window. If increased classification rate is wanted, one can use several convolution windows that operate in parallel. In this case, the combinational clause logic that generates the signal $c_{j}^{b}$ in Fig. 4, would have to be replicated for each convolution window. The outputs from each combinational clause logic instance would have to be input to the OR-gate labeled A in Fig. 4.
@@ IV-E Inference
The inference operation starts with loading an image into the data buffer, followed by sequential generation of the patches, as described in Subsection IV-C. When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq. (3). All ten class sums are generated in parallel, and each class sum is obtained by adding 128 numbers as shown in Fig. 5. For each adder input, a multiplexer (MUX) is employed, which outputs the clause weight, $w_{i,j}$, if $c_{j}=1$, or zero if $c_{j}=0$. The adders are configured in a reduction-tree, with a three-stage pipeline.
Figure 5: Class sum generation for class $i$. The main input signals are the class weights, $w_{i,j}$, and the Boolean clause values, $c_{j}$, where $j\in\{0,\dots,127\}$.
Finally, an argmax module selects the class with the highest class sum, as the predicted class $\hat{y}$. Fig. 6 shows the argmax module which is implemented by combinational logic, using a submodule from [11] in a reduction-tree structure. In Fig. 6, this submodule is shown in the upper right part. The inputs are two class sums, $v_{0}$ and $v_{1}$, and their corresponding 4-bit class labels, label0 and label1. If $v_{1}>v_{0}$, the module feeds forward $v_{1}$ as the $v_{max}$ result, and label1 is output as the maximum argument, $a_{max}$. Otherwise, $v_{0}$ and label0 are selected.
Figure 6: The argmax module, that selects the label corresponding to the largest class sum. Based on a submodule circuit from [11].
Fig. 7 shows a simplified state diagram of the accelerator, and pseudo-code for the inference procedure is described in Algorithm 1.
Algorithm 1 ConvCoTM Accelerator Inference
1:procedure Classify($A$) $\triangleright$ $A$ is a single image example fed to the accelerator by the system processor.
2:    $\triangleright$ From here the algorithm is performed in its entirety by the accelerator.
3:    Reset clause output register.$\triangleright$ Resets the DFF in each instance of the clause logic, ref. Fig. 4.
4:    for $i=0$ to $B-1$ do $\triangleright$ Process $B$ patches, ref. Subsections III-C and IV-C.
5:       Perform in parallel during a single clock cycle:
6:       * Generate patch($i$) of $A$
7:       * Evaluate all clauses, $c_{j}$, $j\in\{0,\dots,n-1\}$, for patch($i$)
8:       * OR the clause output register $c_{j}$ with the combinational clause output $c_{j}^{b}$
9:           and update clause output register $\triangleright$ Ref. Eq. (6) and Fig. 4.
10:    end for
11:    Calculate class sums: $v_{0},v_{1},\dots,v_{m-1}$triangleright Ref. Eq. (3). A tree structure of adders, pipelined in 3 stages, is used per class sum.
12:    return argmax{$v_{0},v_{1},\dots,v_{m-1}$} $\triangleright$ Returns predicted class. Ref. Eq. (4).
13:end procedure
@@
When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C and Fig. 8. This increases the throughput, as image samples can be processed every 372’th clock cycle. Any timing overhead in the system processor will add to the total latency.
Figure 8: Simplified accelerator timing diagram.
@@ V Implementation and Measurement Results
The ConvCoTM accelerator was designed in VHDL and implemented in a 65 nm low-leakage CMOS technology from UMC. We used the Cadence IC design framework for the chip implementation, and the VHDL simulations were performed with Xcelium. Synthesis and Place & Route were performed with Genus and Innovus respectively. To reduce leakage current and to optimize the power consumption for low operating frequencies, relaxed time-constraints were applied, and digital cells with high drive strengths were excluded.
For voltage-drop analysis we applied Voltus, and Conformal was used for logic equivalence checking. For timing verification of the routed design, we used mixed VHDL/Verilog gate-level simulations with back-annotated timing information. Complete gate-level simulations of the accelerator with model loading and inference of the full MNIST test dataset were performed. Design-rule check (DRC) and layer versus schematic (LVS) were performed with Calibre, while GDSII file generation was performed with Virtuoso. The VHDL source code for the ASIC implementation (used for initial FPGA verification) is publicly available at a GitHub repository [40].
A JLCC44 ceramic package housed the chip. For the package pins, there were 5 ground connections, 2 supply voltage connections (3.3 V) for the digital inputs/outputs (IOs), 2 supply connections (1.2 V) for the core cells, and 29 IOs. Six pins were not connected. The number of pins could have been reduced significantly, but as the ASIC was a test chip, we prioritized testability and observability. Fig. 9 shows a photograph and a layout plot of the ASIC. The accelerator’s core area is approximately 2.7 mm2. For the testing and characterization of the ASIC, a printed circuit board (PCB) was developed. The ASIC PCB was connected to a Digilent Zybo Z7-20 development board, equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig. 10. One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC PCB. The processor also read back accelerator status signals, and the predicted class per sample. To interface with the FPGA board, the ASIC’s IOs operated at 3.3 V supply voltage. The accelerator core operated from 0.82 V to 1.2 V.
@@
TABLE II: The ConvCoTM accelerator ASIC’s main characteristics and performance